# Client-Aided Secure Two-Party Computation of Dynamic Controllers

### v2

@@ 5.2 Controller Computation Protocol
1:Controller parameters $A\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times n}$, $B\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times p}$, $C\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times n}$, $D\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times p}$, initial state $x_{0}\in\mathbb{Q}_{\langle k,\ell\rangle}^{n}$, plant output $\hat{y}(t)\in\mathbb{R}^{n}$, bit lengths $k,\ell\in\mathbb{N}$, and modulus $q$
@@
$\ell\geq k-\ell+\log_{2}\frac{c\epsilon^{-1}}{1-\gamma}\quantity(mp\sqrt{n}+n\sqrt{m}),$ (16)
@@
Here, $k$-$\ell$ in (16) represents the bit length of the integer part of the controller parameters and initial state. This result suggests the following procedure for selecting the parameters used in the proposed controller computation protocol: 1) Given the design parameter $\epsilon>0$ and the security parameter $\lambda\in\mathbb{N}$. 2) Fix $k$-$\ell\in\mathbb{N}$ to be sufficiently large. 3) Choose the bit length $\ell\in\mathbb{N}$. 4) Design the controller (2) to satisfy Assumption 1 and Assumption 2. 5) When $D\neq\mathbf{0}_{m\times p}$, check the condition (15). Otherwise, check the condition (16). If $\ell$ does not satisfy the condition, increase $\ell$ and return to step 4). 6) Choose the modulus $q$ to satisfy (14).
@@
During the execution of Protocol 3, each party $P_{i}$ receives the $i$th shares in (9), (10), and (11) from the client and can access public parameters $k$, $\ell$, $q$, and $\lambda$. It also obtains the masked messages $(d,e)$ and $m_{r}$ (if $i=1$) for each invocation of $\mathsf{Mult}$ and $\mathsf{Trunc}$, respectively. Here, the subprotocols are sequentially invoked in the online phase. In this case, the modular composition theorem in [29] guarantees that the security of the entire protocol is equivalent to the weakest security of the subprotocols. Therefore, the proposed protocol achieves statistical security with respect to $\lambda$.
@@ 5.3 Comparison to Previous Works
We compare the theoretical costs in the online phase of the proposed protocol with conventional single-server schemes based on learning with errors (LWE) encryption [14] and ring learning with errors (RLWE) encryption [15, 16]. Table 1 lists their computational complexities on the client and server sides, as well as communication complexities from the client (resp. server) to the server (resp. client) and between the servers, which are detailed below. For a fair comparison, the moduli of the ciphertext spaces and the dimensions of the secret keys in the encryption schemes were set to $q$ and $N$, respectively.
@@
5.3.1 LWE-based encrypted controller in [14]
The authors of [14] represented the controller (2) in the form
@@ 5.3.1 LWE-based encrypted controller in [14]
with $R\in\mathbb{R}^{n\times m}$, $T\in\mathbb{R}^{n\times n}$, $x^{\prime}(t)=Tx(t)\in\mathbb{R}^{n}$, and $\mu(t)=TRu(t)\in\mathbb{R}^{m}$ such that $T(A-RC)T^{-1}\in\mathbb{Z}^{n\times n}$. They encrypted each element of its parameters and signals using the Gentry-Sahai-Waters (GSW) [8] and LWE [5] encryption schemes, respectively. The ciphertexts of the GSW and LWE encryptions are a matrix in $\mathbb{Z}_{q}^{(N+1)\times\zeta(N+1)}$ and a vector in $\mathbb{Z}_{q}^{N+1}$, where $\zeta=\lfloor\log_{\nu}q\rfloor$ for some $\nu\ll q$.
@@
On the server, the encrypted controller requires $n(n+2m+p)+mp$ multiplications between the GSW and LWE ciphertexts and n(n+2m+p4)+m(p1)n(n+2m+p-4)+m(p-1) additions of the LWE ciphertexts. According to [16], each GSW-LWE multiplication comprises $\zeta(N+1)^{2}$ scalar multiplications. In addition, each LWE addition is an $(N+1)$-dimensional vector addition. Consequently, the computational complexity of the server is $O((n+m)(n+p)\zeta N^{2})$.
5.3.2 RLWE-based encrypted controller in [15]
@@
The authors of [15] used the controller representation $u(t)=Hd(t)$ with $d(t)=[y(t-n)^{\top}\ \cdots\ y(t)^{\top}\ u(t-n)^{\top}\ \cdots\ u(t-1)^{\top}]^{\top}$ and encrypted $H\in\mathbb{R}^{m\times(n(m+p)+p)}$ and $d(t)\in\mathbb{R}^{n(m+p)+p}$ using the RLWE encryption [7]. The ciphertext of the RLWE encryption is a pair of polynomials in $R_{q}^{2}$, where $R_{q}\coloneqq\mathbb{Z}_{q}[X]/(X^{N}+1)$ is a polynomial ring consisting of $N$-dimensional polynomials in $X$ over $\mathbb{Z}_{q}$.
@@
Communication complexity: Similar to the LWE-based scheme, for each time $t\in\mathbb{N}_{0}$, the client sends the encryption of $y(t)$ and re-encryption of $u(t)$ and receives the encryption of $u(t)$. Meanwhile, the RLWE-based scheme applied message packing [30] when encrypting $y(t)$ and $u(t)$, which enables the encryption of a vector as a single ciphertext in $R_{q}^{2}$. Therefore, the communication complexity from the client (resp. server) to the server (resp. client) is $O(N\log q)$.
@@
Computational complexity: The message packing and unpacking are respectively performed before encryption and after decryption. The packing and unpacking are implemented by the (inverse) number theoretic transform [15], of which complexity is $O(N\log N)$ [31]. The encryption and decryption include a multiplication of polynomials in $R_{q}$. The $N$-dimensional polynomial multiplication can also be computed using the number theoretic transform. Therefore, the computational complexity of the client is $O(N\log N)$.
@@
The encrypted controller requires $n+1$ multiplications with linearization, $n+m+p+2$ additions, and m+p1m+p-1 rotations of the RLWE ciphertexts. Note that the multiplication in the RLWE encryption increases the ciphertext size from $R_{q}^{2}$ to $R_{q}^{3}$. The linearization reduces it to $R_{q}^{2}$, and the rotation replaces the position of vector elements packed in a ciphertext. Each multiplication and addition process includes four multiplications and two additions in $N$-dimensional polynomials, respectively. The linearization and rotation carry out $2\zeta$ NN-dimensional polynomial multiplications [16], where $\zeta$ is as in the LWE-based scheme. Therefore, the computational complexity of the server is $O((n+m+p)\zeta N\log N)$.
5.3.3 RLWE-based encrypted controller in [16]
@@
When $D=\mathbf{0}_{m\times p}$, the authors of [16] employed a variant of the controller representation in [15], $u(t)=H^{\prime}d^{\prime}(t)$, $d^{\prime}(t)=[y(t-1)^{\top}\ \cdots\ y(t-n)^{\top}\ u(t-1)^{\top}\ \cdots\ u(t-n)^{\top}]^{\top}$, and encrypted $H^{\prime}\in\mathbb{R}^{m\times n(m+p)}$ and $d^{\prime}(t)\in\mathbb{R}^{n(m+p)}$ using the RLWE encryption [6].
Communication complexity: The communication complexities are the same as those in [15].
@@
Computational complexity: This RLWE-based scheme considers a different encoding of data and replaces the addition, linearization, and rotation on the server in [15] with aggregation on the client, which consists of $m^{2}$ scalar additions. This changes the computational complexities of the client and server to $O(m^{2}+N\log N)$ and $O(nN\log N)$, respectively.
@@ 5.3.4 Proposed protocol (Protocol 3)
Compared to the previous schemes, the proposed protocol significantly improves the computational efforts of both the client and servers. Note that the dimension parameter $N$ is typically chosen to be larger than $2^{10}$ [32]. On the other hand, although the proposed protocol reduces the communication complexity from the servers to the client, the reduction in the complexity from the client to the servers is limited to satisfying $(n+m)(n+p)\ll N$. This implies that the proposed protocol incurs high communication costs for large-scale controllers.
@@ 6.1 Brunovsky normal form
It is expected that sparse controller parameters will reduce the number of multiplications of shares by ignoring their zero elements. To transform the parameters of (2) to sparse parameters, we employ the Brunovsky normal form [33].
Definition 6.12 (Brunovsky normal form [34]).
@@ 6.2 Pseudorandom function
By definition, the shares for the party $P_{1}$ are random numbers in $\mathbb{Z}_{q}$ and independent of messages. This characteristic allows us to utilize a pseudorandom function further to reduce the communication cost between the client and parties. A pseudorandom function is a polynomial-time deterministic algorithm $\mathsf{PRF}:\{0,1\}^{\ell_{\mathrm{key}}(\lambda)}\times\{0,1\}^{\ell_{\mathrm{in}}(\lambda)}\to\{0,1\}^{\ell_{\mathrm{out}}(\lambda)}$, which takes an $\ell_{\mathrm{key}}(\lambda)$-bit key and an $\ell_{\mathrm{in}}(\lambda)$-bit string as input and outputs an $\ell_{\mathrm{out}}(\lambda)$-bit string, where $\lambda$ is a security parameter [28]. It is necessary that any polynomial-time algorithm without knowledge of the key cannot distinguish the output from a uniformly random string over $\{0,1\}^{\ell_{\mathrm{out}}(\lambda)}$.
@@ 6.3 Pseudorandom correlation function
Instead of applying the pseudorandom functions, one can use a pseudorandom correlation function [36], which is analogous to a pseudorandom function that generates correlated randomness. A pseudorandom correlation function $\mathsf{PCF}$ is a two-party protocol in which each party $P_{i}$, $i\in\{1,2\}$ is given a key $K_{i}$ and, for some input string $x$, the parties can locally compute correlated strings $(\mathsf{PCF}_{K_{1}}(x),\mathsf{PCF}_{K_{2}}(x))$, where the outputs and random samples from the target correlation are indistinguishable [37].
@@ 7 Numerical Examples
Consider a PID controller of parallel form in [14]. The controller parameters of (2) are $A=\quantity[\begin{smallmatrix}2-N_{d}&N_{d}-1\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, $C=[c_{1}\ c_{2}]$, and $D=d$, where $c_{1}=K_{i}T_{s}-N_{d}^{2}K_{d}/T_{s}$, $c_{2}=N_{d}K_{i}T_{s}-K_{i}T_{s}+N_{d}^{2}K_{d}/T_{s}$, and $d=K_{p}+N_{d}K_{d}/T_{s}$. Here, $K_{p}$, $K_{i}$, and $K_{d}$ are proportional, integral, and derivative gains, respectively, $N_{d}\in\mathbb{N}$ is the parameter of a derivative filter, and $T_{s}\in\mathbb{R}$ is a sampling time. Note that, if $2-N_{d},N_{d}-1\in\mathbb{Z}_{\langle k-\ell\rangle}$, the parameters $A$ and $B$ are already integers. Then, the proposed controller (12) can be simplified as
@@
Suppose that the plant (1) and sampling time are given by the single-input and single-output system in  [38, Eq. (2), with $\alpha=0.2$] and $T_{s}=0.1$ s, respectively. Let $\epsilon=2^{-10}$, $\lambda=80$ bit, $k-\ell=8$ bit, $x_{p,0}=[100\ 100\ 100\ 100]^{\top}$, and $x_{0}=[0\ 0]^{\top}$. For instance, the PID controller with the parameters $A=\quantity[\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, C=[2.73689272.96540833]C=[2.7368927\ {-2.96540833}], and D=5.01071167D=-5.01071167 satisfies Assumption 1, Assumption 2, and (15) with $\ell=32,40,48,56$ bit. The bit length of modulus $q$ can be chosen as $256$ bit to satisfy (14) for all the choices of $\ell$. Fig. 3 shows the input errors $\norm{u(t)-\hat{u}(t)}$. In this figure, the input errors remain below $\epsilon=2^{-10}$ for all $\ell=32,40,58,64$ and $t=0,\dots,50$, thereby supporting Theorem 5.5. Furthermore, the figure suggests that increasing the bit length $\ell$ can reduce the errors.
@@
Suppose the plant is given as the four-tank process in [39] with $h_{1}^{0}=12.4$ cm, $h_{2}^{0}=12.7$ cm, $h_{3}^{0}=1.8$ cm, $h_{4}^{0}=1.4$ cm, $v_{1}^{0}=3$ V, $v_{2}^{0}=3$ V, $\gamma_{1}=0.7$, $\gamma_{2}=0.6$ and $T_{s}=0.5$ s. Let $\epsilon$, $\lambda$, $k$-$\ell$, and $\ell$ be as in the PID control example. Let $x_{p,0}=[10\ 10\ 10\ 10]^{\top}$ and $x_{0}=[0\ 0\ 0\ 0]^{\top}$. The controller with the parameters,

### v3

@@ 5.2 Controller Computation Protocol
1:Controller parameters $A\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times n}$, $B\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times p}$, $C\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times n}$, $D\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times p}$, initial state $x_{0}\in\mathbb{Q}_{\langle k,\ell\rangle}^{n}$, plant output $\hat{y}(t)\in\mathbb{R}^{p}$, bit lengths $k,\ell\in\mathbb{N}$, and modulus $q$
@@
$\ell\geq 2(k-\ell)+\log_{2}\frac{c\epsilon^{-1}}{1-\gamma}(p+1)n\sqrt{m},$ (16)
@@
Here, $k$-$\ell$ in (16) represents the bit length of the integer part of the controller parameters and initial state. This result suggests the following procedure for selecting the parameters used in the proposed controller computation protocol: 1) Given the design parameter $\epsilon>0$ and the security parameter $\lambda\in\mathbb{N}$. 2) Fix $k$-$\ell\in\mathbb{N}$ to be sufficiently large. 3) Choose the bit length $\ell\in\mathbb{N}$. 4) Design the controller (2) to satisfy Assumption 1 and Assumption 2. 5) Check the condition (15) (or (16) when $D=\mathbf{0}_{m\times p}$). If $\ell$ does not satisfy the condition, increase $\ell$ and return to 4). 6) Choose the modulus $q$ to satisfy (14).
@@
During the execution of Protocol 3, each party $P_{i}$ receives the $i$th shares in (9), (10), and (11) from the client and can access public parameters $k$, $\ell$, $q$, and $\lambda$. It also obtains the masked messages $(d,e)$ and $m_{r}$ (if $i=1$) for each invocation of $\mathsf{Mult}$ and $\mathsf{Trunc}$, respectively. Here, the subprotocols are sequentially invoked in the online phase. In this case, the modular composition theorem in [Canetti2000-cl] guarantees that the security of the entire protocol is equivalent to the weakest security of the subprotocols. Therefore, the proposed protocol achieves statistical security with respect to $\lambda$.
@@ 5.3 Comparison to Previous Works
We compare the theoretical costs in the online phase of the proposed protocol with conventional single-server schemes based on learning with errors (LWE) encryption [Kim2023-pk] and ring learning with errors (RLWE) encryption [Teranishi2024-ha, Lee2025-jo]. Table 1 lists their computational complexities on the client and server sides, as well as communication complexities from the client (resp. server) to the server (resp. client) and between the servers, which are detailed below. For a fair comparison, the moduli of the ciphertext spaces and the dimensions of the secret keys in the encryption schemes were set to $q$ and $N$, respectively.
@@
5.3.1 LWE-based encrypted controller in [Kim2023-pk]
The authors of [Kim2023-pk] represented the controller (2) in the form
@@ 5.3.1 LWE-based encrypted controller in [Kim2023-pk]
with $R\in\mathbb{R}^{n\times m}$, $T\in\mathbb{R}^{n\times n}$, $x^{\prime}(t)=Tx(t)\in\mathbb{R}^{n}$, and $\mu(t)=TRu(t)\in\mathbb{R}^{m}$ such that $T(A-RC)T^{-1}\in\mathbb{Z}^{n\times n}$. They encrypted each element of its parameters and signals using the Gentry-Sahai-Waters (GSW) [Gentry2013-eq] and LWE [Regev2009-ys] encryption schemes, respectively. The ciphertexts of the GSW and LWE encryptions are a matrix in $\mathbb{Z}_{q}^{(N+1)\times\zeta(N+1)}$ and a vector in $\mathbb{Z}_{q}^{N+1}$, where $\zeta=\lfloor\log_{\nu}q\rfloor$ for some $\nu\ll q$.
@@
On the server, the encrypted controller requires $n(n+2m+p)+mp$ multiplications between the GSW and LWE ciphertexts and n(n+2m+p4)+m(p1)n(n+2m+p-4)+m(p-1) additions of the LWE ciphertexts. According to [Lee2025-jo], each GSW-LWE multiplication comprises $\zeta(N+1)^{2}$ scalar multiplications. In addition, each LWE addition is an $(N+1)$-dimensional vector addition. Consequently, the computational complexity of the server is $O((n+m)(n+p)\zeta N^{2})$.
5.3.2 RLWE-based encrypted controller in [Teranishi2024-ha]
@@
The authors of [Teranishi2024-ha] used the controller representation $u(t)=Hd(t)$ with $d(t)=[y(t-n)^{\top}\ \cdots\ y(t)^{\top}\ u(t-n)^{\top}\ \cdots\ u(t-1)^{\top}]^{\top}$ and encrypted $H\in\mathbb{R}^{m\times(n(m+p)+p)}$ and $d(t)\in\mathbb{R}^{n(m+p)+p}$ using the RLWE encryption [Fan2012-jt]. The ciphertext of the RLWE encryption is a pair of polynomials in $R_{q}^{2}$, where $R_{q}\coloneqq\mathbb{Z}_{q}[X]/(X^{N}+1)$ is a polynomial ring consisting of $N$-dimensional polynomials in $X$ over $\mathbb{Z}_{q}$.
@@
Communication complexity: Similar to the LWE-based scheme, for each time $t\in\mathbb{N}_{0}$, the client sends the encryption of $y(t)$ and re-encryption of $u(t)$ and receives the encryption of $u(t)$. Meanwhile, the RLWE-based scheme applied message packing [Smart2012-pc] when encrypting $y(t)$ and $u(t)$, which enables the encryption of a vector as a single ciphertext in $R_{q}^{2}$. Therefore, the communication complexity from the client (resp. server) to the server (resp. client) is $O(N\log q)$.
@@
Computational complexity: The message packing and unpacking are respectively performed before encryption and after decryption. The packing and unpacking are implemented by the (inverse) number theoretic transform [Teranishi2024-ha], of which complexity is $O(N\log N)$ [Satriawan2023-gz]. The encryption and decryption include a multiplication of polynomials in $R_{q}$. The $N$-dimensional polynomial multiplication can also be computed using the number theoretic transform. Therefore, the computational complexity of the client is $O(N\log N)$.
@@
The encrypted controller requires $n+1$ multiplications with linearization, $n+m+p+2$ additions, and m+p1m+p-1 rotations of the RLWE ciphertexts. Note that the multiplication in the RLWE encryption increases the ciphertext size from $R_{q}^{2}$ to $R_{q}^{3}$. The linearization reduces it to $R_{q}^{2}$, and the rotation replaces the position of vector elements packed in a ciphertext. Each multiplication and addition process includes four multiplications and two additions in $N$-dimensional polynomials, respectively. The linearization and rotation carry out $2\zeta$ NN-dimensional polynomial multiplications [Lee2025-jo], where $\zeta$ is as in the LWE-based scheme. Therefore, the computational complexity of the server is $O((n+m+p)\zeta N\log N)$.
5.3.3 RLWE-based encrypted controller in [Lee2025-jo]
@@
When $D=\mathbf{0}_{m\times p}$, the authors of [Lee2025-jo] employed a variant of the controller representation in [Teranishi2024-ha], $u(t)=H^{\prime}d^{\prime}(t)$, $d^{\prime}(t)=[y(t-1)^{\top}\ \cdots\ y(t-n)^{\top}\ u(t-1)^{\top}\ \cdots\ u(t-n)^{\top}]^{\top}$, and encrypted $H^{\prime}\in\mathbb{R}^{m\times n(m+p)}$ and $d^{\prime}(t)\in\mathbb{R}^{n(m+p)}$ using the RLWE encryption [Brakerski2014-ny].
Communication complexity: The communication complexities are the same as those in [Teranishi2024-ha].
@@
Computational complexity: This RLWE-based scheme considers a different encoding of data and replaces the addition, linearization, and rotation on the server in [Teranishi2024-ha] with aggregation on the client, which consists of $m^{2}$ scalar additions. This changes the computational complexities of the client and server to $O(m^{2}+N\log N)$ and $O(nN\log N)$, respectively.
@@ 5.3.4 Proposed protocol (Protocol 3)
Compared to the previous schemes, the proposed protocol significantly improves the computational efforts of both the client and servers. Note that the dimension parameter $N$ is typically chosen to be larger than $2^{10}$ [HES2018]. On the other hand, although the proposed protocol reduces the communication complexity from the servers to the client, the reduction in the complexity from the client to the servers is limited to satisfying $(n+m)(n+p)\ll N$. This implies that the proposed protocol incurs high communication costs for large-scale controllers.
@@ 6.1 Brunovsky normal form
It is expected that sparse controller parameters will reduce the number of multiplications of shares by ignoring their zero elements. To transform the parameters of (2) to sparse parameters, we employ the Brunovsky normal form [Brunovsk-`y1970-kk].
Definition 6.12 (Brunovsky normal form [Tabuada2003-fp]).
@@ 6.2 Pseudorandom function
By definition, the shares for the party $P_{1}$ are random numbers in $\mathbb{Z}_{q}$ and independent of messages. This characteristic allows us to utilize a pseudorandom function further to reduce the communication cost between the client and parties. A pseudorandom function is a polynomial-time deterministic algorithm $\mathsf{PRF}:\{0,1\}^{\ell_{\mathrm{key}}(\lambda)}\times\{0,1\}^{\ell_{\mathrm{in}}(\lambda)}\to\{0,1\}^{\ell_{\mathrm{out}}(\lambda)}$, which takes an $\ell_{\mathrm{key}}(\lambda)$-bit key and an $\ell_{\mathrm{in}}(\lambda)$-bit string as input and outputs an $\ell_{\mathrm{out}}(\lambda)$-bit string, where $\lambda$ is a security parameter [Katz2014-kb]. It is necessary that any polynomial-time algorithm without knowledge of the key cannot distinguish the output from a uniformly random string over $\{0,1\}^{\ell_{\mathrm{out}}(\lambda)}$.
@@ 6.3 Pseudorandom correlation function
Instead of applying the pseudorandom functions, one can use a pseudorandom correlation function [Boyle2020-sh], which is analogous to a pseudorandom function that generates correlated randomness. A pseudorandom correlation function $\mathsf{PCF}$ is a two-party protocol in which each party $P_{i}$, $i\in\{1,2\}$ is given a key $K_{i}$ and, for some input string $x$, the parties can locally compute correlated strings $(\mathsf{PCF}_{K_{1}}(x),\mathsf{PCF}_{K_{2}}(x))$, where the outputs and random samples from the target correlation are indistinguishable [Couteau2023-fs].
@@ 7 Numerical Examples
Consider a PID controller of parallel form in [Kim2023-pk]. The controller parameters of (2) are $A=\quantity[\begin{smallmatrix}2-N_{d}&N_{d}-1\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, $C=[c_{1}\ c_{2}]$, and $D=d$, where $c_{1}=K_{i}T_{s}-N_{d}^{2}K_{d}/T_{s}$, $c_{2}=N_{d}K_{i}T_{s}-K_{i}T_{s}+N_{d}^{2}K_{d}/T_{s}$, and $d=K_{p}+N_{d}K_{d}/T_{s}$. Here, $K_{p}$, $K_{i}$, and $K_{d}$ are proportional, integral, and derivative gains, respectively, $N_{d}\in\mathbb{N}$ is the parameter of a derivative filter, and $T_{s}\in\mathbb{R}$ is a sampling time. Note that, if $2-N_{d},N_{d}-1\in\mathbb{Z}_{\langle k-\ell\rangle}$, the parameters $A$ and $B$ are already integers. Then, the proposed controller (12) can be simplified as
@@
Suppose that the plant (1) and sampling time are given by the single-input and single-output system in  [Astrom2000-ta, Eq. (2), with $\alpha=0.2$] and $T_{s}=0.1$ s, respectively. Let $\epsilon=2^{-10}$, $\lambda=80$ bit, $k-\ell=8$ bit, $x_{p,0}=[100\ 100\ 100\ 100]^{\top}$, and $x_{0}=[0\ 0]^{\top}$. For instance, the PID controller with the parameters $A=\quantity[\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, C=[2.73689272.96540833]C=[2.7368927\ {-2.96540833}], and D=5.01071167D=-5.01071167 satisfies Assumption 1, Assumption 2, and (15) with $\ell=32,40,48,56$ bit. The bit length of modulus $q$ can be chosen as $256$ bit to satisfy (14) for all the choices of $\ell$. Fig. 3 shows the input errors $\norm{u(t)-\hat{u}(t)}$. In this figure, the input errors remain below $\epsilon=2^{-10}$ for all $\ell=32,40,58,64$ and $t=0,\dots,50$, thereby supporting Theorem 5.5. Furthermore, the figure suggests that increasing the bit length $\ell$ can reduce the errors.
@@
Suppose the plant is given as the four-tank process in [Johansson2000-il] with $h_{1}^{0}=12.4$ cm, $h_{2}^{0}=12.7$ cm, $h_{3}^{0}=1.8$ cm, $h_{4}^{0}=1.4$ cm, $v_{1}^{0}=3$ V, $v_{2}^{0}=3$ V, $\gamma_{1}=0.7$, $\gamma_{2}=0.6$ and $T_{s}=0.5$ s. Let $\epsilon$, $\lambda$, $k$-$\ell$, and $\ell$ be as in the PID control example. Let $x_{p,0}=[10\ 10\ 10\ 10]^{\top}$ and $x_{0}=[0\ 0\ 0\ 0]^{\top}$. The controller with the parameters,